# Revisiting the propagation of highly-energetic gamma rays in the Galaxy

### v1

@@ Acknowledgments
The simulation data performed and employed in this work are accessible for downloading at the data repository of the DAMASCO group website. The files contain also the EM cascade electrons reaching the observer, despite they are not used in our analysis. We encourage those interested in the data to reach out the authors for discussions about data structuring and analysis.
@@ Appendix A Magnetic field along the line of sight
The GMF structure in JF12+ comprehends the regular, striated, and turbulent components, as explained in Sec. 2.2. In Fig. 8 and Fig. 9, two GMF configurations are depicted along two lines of sight, namely the one related to the $S_{<}$ source, $\sim 6\;\text{kpc}$ distant, and the other to $S_{1}$ source. While in the former the line of sight goes through many GMF spirals, the latter crosses only one spiral. This fact is reflected in the interplay between regular and turbulent components. In the ratios we neglect the contribution from the random striated field since it is expected to be negligible in most of the accounted galactic regions, according to JF12+.
@@ Appendix B Observer size
In general, the purpose in choosing a certain radius is to find a compromise between the smallest observer size and statistics, thus the amount of detected gamma rays. The latter is directly related to computational time. To avoid altering the results by systematic trends due to a specific choice, it is important to evaluate two relevant quantities: the electrons mean free paths and the typical length scales of the curvature of magnetic field lines. Both quantities are due to the environmental properties in the surroundings of the observer. The former is the inverse of Eq. 2.1 in the case of IC scattering, reported in Fig. 2, thus the electron mean free path. The latter is interpreted as the magnetic field curvature length scale; it is expressed as (×$\big{(}\frac{\vec{\nabla}\times\vec{B}}{\vec{B}}\big{)}^{-1}$, where $\vec{B}$ is the magnetic field. Simulations, that include more accurate and detailed ISRF and GMF models in the Earth surroundings, will need to take into account the two lengths, just introduced. In this case, to reduce the impact of systematics on simulation results, the observer radius will need to be chosen as the minimum between the electron interaction mean distances and the magnetic field curvature scale.
In Fig. 11, we show the observed energy spectra in dependence to the spherical observer radius. It ranges from the geometrical one ($\sim$ 110 pc) to the average magnetic field curvature length near the Earth ($\sim$ 45 pc) for the JF12+ model. The energy spectra features, discussed in Sec. 3.1, are similar for the six observer radii set in the simulation.
@@ Appendix C Arrival direction maps
This appendix contains the spatial count maps for the ten sources simulated in this work, as discussed in Sec. 3.2. The count maps of $S_{\lessgtr}$ sources are in Fig. 12 and Fig. 13, respectively for the closest and farthest sources. The ones corresponding to $S_{1}$ and $S_{2}$ are shown in Fig. 15 and Fig. 15, respectively. The amount of injected gamma rays in the simulation of each source is $10^{7}$. To note that the counts on the colorbars are not effective since, in our simulations, the thinning procedure is applied to sample the secondary particles in the cascade [74].

### v4

@@ Acknowledgements.
The simulation data performed and employed in this work are accessible for downloading at the data repository of the DAMASCO group website. The files contain also the EM cascade electrons reaching the observer, despite they are not accounted in our analysis. We encourage those interested in the data to reach out the authors for discussions about data structuring and analysis.
@@ Appendix A Magnetic field along the line of sight
The GMF structure in JF12+ comprehends the regular, striated, and turbulent components, as explained in Sec. II.2. In Fig. 14 and Fig. 15, two GMF configurations are depicted along two lines of sight, namely the one related to the $S_{<}$ source, $\sim 6\;\text{kpc}$ distant, and the other to $S_{1}$ source. While in the former the line of sight goes through many GMF spirals, the latter crosses only one spiral. This fact is reflected in the interplay between regular and turbulent components. In the ratios we neglect the contribution from the random striated field since it is expected to be negligible in most of the accounted galactic regions, according to JF12+.
@@ Appendix B Observer size
In general, the purpose in choosing a certain radius is to find a compromise between the smallest observer size and statistics, thus the amount of detected gamma rays. The latter is directly related to computational time. To avoid altering the results by systematic trends due to a specific choice, it is important to evaluate two relevant quantities: the electrons mean free paths and the typical length scales of the curvature of magnetic field lines. Both quantities are due to the environmental properties in the surroundings of the observer. The former is the inverse of Eq. 1 in the case of IC scattering, reported in Fig. 2, thus the electron mean free path. The latter is interpreted as the magnetic field curvature length scale; it is expressed as (×$\big{(}\frac{\vec{\nabla}\times\vec{B}}{\vec{B}}\big{)}^{-1}$, where $\vec{B}$ is the magnetic field. Simulations, that include more accurate and detailed ISRF and GMF models in the Earth surroundings, will need to take into account the two lengths, just introduced. In this case, to reduce the impact of systematics on simulation results, the observer radius will need to be chosen as the minimum between the electron interaction mean distances and the magnetic field curvature scale.
In Fig. 17, we show the observed energy spectra in dependence to the spherical observer radius. It ranges from the geometrical one ($\sim$ 110 pc) to the average magnetic field curvature length near the Earth ($\sim$ 45 pc) for the JF12+ model. The energy spectra features, discussed in Sec. III.1, are similar for the six observer radii set in the simulation.
@@ Appendix C Arrival direction maps
This appendix contains the spatial count maps for the ten sources simulated in this work, as discussed in Sec. III.2. The count maps of $S_{\lessgtr}$ sources are in Fig. 22 and Fig. 27, respectively for the closest and farthest sources. The ones corresponding to $S_{1}$ and $S_{2}$ are shown in Fig. 29 and Fig. 29, respectively. The amount of injected gamma rays in the simulation of each source is $10^{7}$. To note that the counts on the colorbars are not effective since, in our simulations, the thinning procedure is applied to sample the secondary particles in the cascade [77].
Figure 18: *